Sol Summary – Oct 31st


Summary Title: Halloween Heliograph
Author’s name: Guy Murphy

Mission Status: Active.

Sol Activity Summary: EVA 4 was undertaken this morning to the chosen sample collection site for the micrometeorite study. During the EVA a heliograph was experimentally used to signal the hab from the field. Baseline micro/macro nutrients have been inputted into the food stock database. We completed tying the remaining untied tunnel tarpaulin eye loops to all the tunnels. 3 cooked meals prepared, including a tuna risotto with pesto sauce for dinner with an apple crumble desert. General hab housework duties were undertaken.

Look Ahead Plan: The EVA program will continue for the remainder of the week subject to favourable weather conditions.

Anomalies in work: None.

Weather: After a very cold night, it was another clear, sunny day with no precipitation.

Crew Physical Status: Crew in good health.

EVA: Successfully undertaken, see EVA report.

Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.

Support Requested: We are still awaiting the delivery of food supplies, the absence of which is beginning to constrain our diet. Several plastic funnels to assist with the reuse of cooking water and several tea towels are also requested.

Copyright © The Mars Society. All rights reserved. | Main Site